# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 4. Koszul duality
So we need to see that the unit and counit of the adjunction of 2. and of the adjunction (11) are equivalences. We will show that the unit $\eta$ of the adjunction of 2. is an equivalence. The case of the counit of the adjunction of 2. and the other cases are similar.
We will prove that for every ${\mathrm{k}}\geq 1$ the following statement ($\ast$) holds: for every ${\mathrm{n}}\geq 1$ and every right ${\mathcal{O}}$-module ${\mathrm{X}}$ that vanishes below degree ${\mathrm{n}}$, i.e. that belongs to ${\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}),$ the following morphism is an equivalence:
$(\eta_{\mathrm{X}})_{\mathrm{k}}:{\mathrm{X}}_{\mathrm{k}}\to(({\mathrm{X}}\circ_{\mathcal{O}}{\mathrm{triv}})\circ^{{\mathcal{O}}^{\vee}}{\mathrm{triv}})_{\mathrm{k}}.$
By Lemma 2.39 the objects ${\mathrm{X}}\circ_{\mathcal{O}}{\mathrm{triv}}$ and so $({\mathrm{X}}\circ_{\mathcal{O}}{\mathrm{triv}})\circ^{{\mathcal{O}}^{\vee}}{\mathrm{triv}}$ belong to ${\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}$. Hence $\eta_{\mathrm{k}}$ is an equivalence if ${\mathrm{n}}>{\mathrm{k}}$. Thus it remains to show ($\ast$) for ${\mathrm{n}}\leq{\mathrm{k}},$ which we prove by descending induction on ${\mathrm{n}}.$ By Lemma 2.43 there is an ${\mathrm{X}}^{\prime}\in{\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}+1})$ and a morphism ${\mathrm{X}}^{\prime}\to{\mathrm{X}}$ in ${\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}})$ that induces an equivalence in degrees larger than ${\mathrm{n}}.$ By Lemma 2.46 the cofiber ${\mathrm{X}}^{\prime\prime}$ of the morphism ${\mathrm{X}}^{\prime}\to{\mathrm{X}}$ in the stable $\infty$-category ${\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})$ is the trivial right ${\mathcal{O}}$-module concentrated in degree ${\mathrm{n}}$ with value ${\mathrm{X}}_{\mathrm{n}}$ since ${\mathcal{O}}_{1}={\mathbb{1}}$. There is a commutative square
@@
of the adjunction of 2. sends cofree right ${\mathcal{O}}^{\vee}$-comodules to trivial right ${\mathcal{O}}$-modules.
On the other hand the left adjoint of adjunction 2. sends trivial right ${\mathcal{O}}$-modules to cofree right ${\mathcal{O}}^{\vee}$-comodules since the composition product on ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ preserves small sifted colimits in each component.
@@
Let ${\mathcal{C}}$ be a stable symmetric monoidal $\infty$-category compatible with small colimits that admits totalizations. Let ${\mathcal{O}}$ be a non-unital $\infty$-operad in ${\mathcal{C}}$ such that ${\mathcal{O}}_{1}={\mathbb{1}}$. For every ${\mathrm{n}}\in{\mathbb{N}}$ there is a canonical equivalence of right ${\mathcal{O}}$-modules:
${\mathrm{triv}}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{O}})}{\mathcal{O}}\simeq\tau_{\mathrm{n}}({\mathcal{O}}^{\vee})\circ^{{\mathcal{O}}^{\vee}}{\mathrm{triv}}.$
@@
Let $\psi:\mathsf{f}_{\mathrm{n}}({\mathcal{O}})\to{\mathcal{O}}$ be the canonical map of $\infty$-operads.
@@
[…]
@@
${\mathrm{triv}}\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}\simeq\mathrm{colim}_{{\mathrm{n}}\geq 1}((\tau_{\mathrm{n}}({\mathcal{Q}})\circ^{{\mathcal{Q}}}{\mathrm{triv}})\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}).$
@@
We first assume that ${\mathcal{C}}$ is a stable presentably symmetric monoidal $\infty$-category. In this case by Remark 2.42 for every ${\mathrm{n}}\geq 1$ the functor $\tau_{\mathrm{n}}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq{\mathrm{n}}}\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}$ admits a left adjoint $\mathsf{f}_{\mathrm{n}}$. By [21, Remark 4.23.] the canonical map of $\infty$-operads $\mathrm{colim}_{{\mathrm{n}}\geq 1}\mathsf{f}_{\mathrm{n}}({\mathcal{Q}}^{\vee})\to{\mathcal{Q}}^{\vee}$ is an equivalence. By Lemma 2.42 this implies that the canonical map
$\mathrm{colim}_{{\mathrm{n}}\geq 1}{\mathcal{Q}}^{\vee}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{Q}}^{\vee})}{\mathrm{A}}\to{\mathrm{A}}$
of ${\mathcal{Q}}^{\vee}$-algebras in ${\mathcal{C}}$ is an equivalence. Hence the canonical map
$\mathrm{colim}_{{\mathrm{n}}\geq 1}{\mathrm{triv}}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{Q}}^{\vee})}{\mathrm{A}}\simeq{\mathrm{triv}}\circ_{{\mathcal{Q}}^{\vee}}(\mathrm{colim}_{{\mathrm{n}}\geq 1}{\mathcal{Q}}^{\vee}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{Q}}^{\vee})}{\mathrm{A}})\to{\mathrm{triv}}\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}$
is an equivalence. By Lemma 4.8 and Proposition 4.6 there is a canonical equivalence ${\mathrm{triv}}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{Q}}^{\vee})}{\mathcal{Q}}^{\vee}\simeq\tau_{\mathrm{n}}({\mathcal{Q}})\circ^{{\mathcal{Q}}}{\mathrm{triv}}$ of right ${\mathcal{Q}}^{\vee}$-modules. So we obtain a canonical equivalence
$\theta_{\mathcal{C}}:(\tau_{\mathrm{n}}({\mathcal{Q}})\circ^{{\mathcal{Q}}}{\mathrm{triv}})\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}\simeq({\mathrm{triv}}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{Q}}^{\vee})}{\mathcal{Q}}^{\vee})\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}\to{\mathrm{triv}}\circ_{\mathsf{f}_{\mathrm{n}}({\mathcal{Q}}^{\vee})}{\mathrm{A}}$
@@
Now let ${\mathcal{C}}$ be an arbitrary stable symmetric monoidal $\infty$-category that admits small colimits and totalizations. By Corollary 7.10 there are symmetric monoidal exact embeddings $\iota:{\mathcal{C}}\subset{\mathcal{C}}^{\prime},{\mathrm{j}}:{\mathcal{C}}^{\prime}\subset{\mathcal{C}}^{\prime\prime}$ into stable symmetric monoidal $\infty$-categories, where ${\mathcal{C}}^{\prime\prime}$ is a presentably symmetric monoidal $\infty$-category, ${\mathcal{C}}^{\prime}$ is compatible with small colimits, the first embedding $\iota:{\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ admits a left adjoint ${\mathrm{L}}$, the second embedding ${\mathrm{j}}:{\mathcal{C}}^{\prime}\subset{\mathcal{C}}^{\prime\prime}$ preserves small colimits and the whole embedding ${\mathrm{j}}\circ\iota:{\mathcal{C}}\subset{\mathcal{C}}^{\prime\prime}$ preserves limits. We obtain induced embeddings
$\iota:{\mathrm{coOp}}({\mathcal{C}})\subset{\mathrm{coOp}}({\mathcal{C}}^{\prime}),{\mathrm{j}}:{\mathrm{coOp}}({\mathcal{C}}^{\prime})\subset{\mathrm{coOp}}({\mathcal{C}}^{\prime\prime}),$
$\iota:\mathrm{Alg}_{{\mathcal{Q}}^{\vee}}({\mathcal{C}})\subset\mathrm{Alg}_{{\mathcal{Q}}^{\vee}}({\mathcal{C}}^{\prime}),{\mathrm{j}}:\mathrm{Alg}_{{\mathcal{Q}}^{\vee}}({\mathcal{C}}^{\prime})\subset\subset\mathrm{Alg}_{{\mathcal{Q}}^{\vee}}({\mathcal{C}}^{\prime\prime})$
@@
${\mathrm{j}}:{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1})\subset{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime\prime})_{\geq 1}),$
$\iota:{\mathrm{coRMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\subset{\mathrm{coRMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}),$
${\mathrm{j}}:{\mathrm{coRMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1})\subset{\mathrm{coRMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime\prime})_{\geq 1}).$
Since the embedding ${\mathrm{j}}\circ\iota:{\mathcal{C}}\subset{\mathcal{C}}^{\prime\prime}$ preserves limits, the functor
$(-)\circ^{{\mathrm{j}}\iota({\mathcal{Q}})}{\mathrm{j}}\iota({\mathrm{triv}})\simeq(-)\ast^{{\mathrm{j}}\iota({\mathcal{Q}})}{\mathrm{j}}\iota({\mathrm{triv}}):{\mathrm{coRMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime\prime})_{\geq 1})\to{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime\prime})_{\geq 1})$
@@
$(-)\circ^{{\mathcal{Q}}}{\mathrm{triv}}\simeq(-)*^{{\mathcal{Q}}}{\mathrm{triv}}:{\mathrm{coRMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\to{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}).$
Let ${\mathrm{T}}:=\tau_{\mathrm{n}}({\mathcal{Q}})\circ^{{\mathcal{Q}}}{\mathrm{triv}}$. Consequently, there is a canonical equivalence:
${\mathrm{j}}\iota({\mathrm{T}})\simeq{\mathrm{j}}\iota(\tau_{\mathrm{n}}({\mathcal{Q}}))\circ^{{\mathrm{j}}\iota({\mathcal{Q}})}{\mathrm{j}}\iota({\mathrm{triv}})\simeq\tau_{\mathrm{n}}({\mathrm{j}}\iota({\mathcal{Q}}))\circ^{{\mathrm{j}}\iota({\mathcal{Q}})}{\mathrm{j}}\iota({\mathrm{triv}}).$
@@
${\mathrm{j}}\iota({\mathrm{triv}})\circ_{{\mathrm{j}}\iota({\mathcal{Q}})^{\vee}}{\mathrm{j}}\iota({\mathrm{A}})\simeq\mathrm{colim}_{{\mathrm{n}}\geq 1}((\tau_{\mathrm{n}}({\mathrm{j}}\iota({\mathcal{Q}}))\circ^{{\mathrm{j}}\iota({\mathcal{Q}})}{\mathrm{j}}\iota({\mathrm{triv}}))\circ_{{\mathrm{j}}\iota({\mathcal{Q}})^{\vee}}{\mathrm{j}}\iota({\mathrm{A}}))$
$\simeq\mathrm{colim}_{{\mathrm{n}}\geq 1}({\mathrm{j}}\iota({\mathrm{T}})\circ_{{\mathrm{j}}\iota({\mathcal{Q}})^{\vee}}{\mathrm{j}}\iota({\mathrm{A}})).$
Since the embedding ${\mathrm{j}}:{\mathcal{C}}^{\prime}\subset{\mathcal{C}}^{\prime\prime}$ preserves small colimits, there are canonical equivalences
${\mathrm{j}}\iota({\mathrm{triv}})\circ_{{\mathrm{j}}\iota({\mathcal{Q}})^{\vee}}{\mathrm{j}}\iota({\mathrm{A}})\simeq{\mathrm{j}}\iota({\mathrm{triv}})\circ_{{\mathrm{j}}\iota({\mathcal{Q}}^{\vee})}{\mathrm{j}}\iota({\mathrm{A}})\simeq{\mathrm{j}}(\iota({\mathrm{triv}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})),$
${\mathrm{j}}\iota({\mathrm{T}})\circ_{{\mathrm{j}}\iota({\mathcal{Q}})^{\vee}}{\mathrm{j}}\iota({\mathrm{A}})\simeq{\mathrm{j}}\iota({\mathrm{T}})\circ_{{\mathrm{j}}\iota({\mathcal{Q}}^{\vee})}{\mathrm{j}}\iota({\mathrm{A}})\simeq{\mathrm{j}}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})).$
@@
${\mathrm{j}}(\iota({\mathrm{triv}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))\simeq\mathrm{colim}_{{\mathrm{n}}\geq 1}{\mathrm{j}}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))$
$\simeq{\mathrm{j}}(\mathrm{colim}_{{\mathrm{n}}\geq 1}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})))$
@@
$\iota({\mathrm{triv}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})\simeq\mathrm{colim}_{{\mathrm{n}}\geq 1}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))$
@@
${\mathrm{triv}}\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}\simeq{\mathrm{L}}(\iota({\mathrm{triv}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))\simeq{\mathrm{L}}(\mathrm{colim}_{{\mathrm{n}}\geq 1}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})))\simeq$
$\mathrm{colim}_{{\mathrm{n}}\geq 1}{\mathrm{L}}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))\simeq\mathrm{colim}_{{\mathrm{n}}\geq 1}({\mathrm{T}}\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}).$
@@
Let ${\mathcal{C}}$ be a stable symmetric monoidal $\infty$-category that admits small limits, ${\mathcal{O}}$ a non-unital $\infty$-operad in ${\mathcal{C}}$ such that ${\mathcal{O}}_{1}={\mathbb{1}}$. For every ${\mathrm{C}}\in{\mathrm{coAlg}}^{\mathrm{dp}}_{{\mathcal{O}}^{\vee}}({\mathcal{C}})$ in ${\mathcal{C}}$ there is a canonical equivalence in ${\mathcal{C}}$:
${\mathrm{triv}}*^{{\mathcal{O}}^{\vee}}{\mathrm{C}}\simeq\lim_{{\mathrm{n}}\geq 1}((\tau_{\mathrm{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})*^{{\mathcal{O}}^{\vee}}{\mathrm{C}}).$

### v2

@@ 3. Koszul duality
So we need to see that the unit and counit of the adjunction of 2. and of the adjunction (3.2) are equivalences. We will show that the unit $\eta$ of the adjunction of 2. is an equivalence. The case of the counit of the adjunction of 2. and the other cases are similar.
We will prove that for every ${k}\geq 1$ the following statement ($\ast$) holds: for every ${n}\geq 1$ and every right $\mathcal{O}$}-module $\mathrm{X}$} that vanishes below degree n{n}, i.e. that belongs to ${\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}),$ the following morphism is an equivalence:
$(\eta_{\mathrm{X}})_{k}:{\mathrm{X}}_{k}\to(({\mathrm{X}}\circ_{\mathcal{O}}{\mathrm{triv}})\circ^{{\mathcal{O}}^{\vee}}{\mathrm{triv}})_{k}.$
By Lemma 2.39 the objects ${\mathrm{X}}\circ_{\mathcal{O}}{\mathrm{triv}}$ and so $({\mathrm{X}}\circ_{\mathcal{O}}{\mathrm{triv}})\circ^{{\mathcal{O}}^{\vee}}{\mathrm{triv}}$ belong to ${\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}$. Hence $\eta_{k}$ is an equivalence if n>k{n}>{k}. Thus it remains to show ($\ast$) for ${n}\leq{k},$ which we prove by descending induction on n.{n}. By Lemma 2.43 there is an ${\mathrm{X}}^{\prime}\in{\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}+1})$ and a morphism ${\mathrm{X}}^{\prime}\to{\mathrm{X}}$ in ${\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}})$ that induces an equivalence in degrees larger than n.{n}. By Lemma 2.46 the cofiber $\mathrm{X}$$^{\prime\prime}$ of the morphism ${\mathrm{X}}^{\prime}\to{\mathrm{X}}$ in the stable $\infty$-category ${\mathrm{RMod}}_{\mathcal{O}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})$ is the trivial right $\mathcal{O}$}-module concentrated in degree n{n} with value ${\mathrm{X}}_{n}$ since ${\mathcal{O}}_{1}={\mathbb{1}}$. There is a commutative square
@@
[…]
@@
Let $\mathcal{C}$} be a stable symmetric monoidal $\infty$-category compatible with small colimits that admits totalizations. Let $\mathcal{O}$} be a non-unital $\infty$-operad in $\mathcal{C}$} such that ${\mathcal{O}}_{1}={\mathbb{1}}$. For every ${n}\in{\mathbb{N}}$ there is a canonical equivalence of right $\mathcal{O}$}-modules:
${\mathrm{triv}}\circ_{{\mathrm{f}}_{n}({\mathcal{O}})}{\mathcal{O}}\simeq\tau_{n}({\mathcal{O}}^{\vee})\circ^{{\mathcal{O}}^{\vee}}{\mathrm{triv}}.$
@@
Let $\psi:{\mathrm{f}}_{n}({\mathcal{O}})\to{\mathcal{O}}$ be the canonical map of $\infty$-operads.
@@
${\mathrm{co}{\mathrm{RMod}}_{{\mathcal{O}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\psi^{\vee})^{\ast}}$${{\mathrm{RMod}}_{{\mathcal{O}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi^{\ast}}$${\mathrm{co}{\mathrm{RMod}}_{\tau_{n}({\mathcal{O}}^{\vee})}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\psi^{\vee})_{\ast}}$${{\mathrm{RMod}}_{{\mathrm{f}}_{n}({\mathcal{O}})}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{\ast}}$${\mathrm{co}{\mathrm{RMod}}_{{\mathcal{O}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{RMod}}_{{\mathcal{O}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}).}$
@@
${\mathrm{triv}}\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}\simeq\mathrm{colim}_{{n}\geq 1}((\tau_{n}({\mathcal{Q}})\circ^{{\mathcal{Q}}}{\mathrm{triv}})\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}).$
@@
We first assume that $\mathcal{C}$} is a stable presentably symmetric monoidal $\infty$-category. In this case by Remark 2.42 for every ${n}\geq 1$ the functor $\tau_{n}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq{n}}\subset{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}$ admits a left adjoint ${\mathrm{f}}_{n}$. By [30, Remark 4.23.] the canonical map of $\infty$-operads $\mathrm{colim}_{{n}\geq 1}{\mathrm{f}}_{n}({\mathcal{Q}}^{\vee})\to{\mathcal{Q}}^{\vee}$ is an equivalence. By Lemma 2.42 this implies that the canonical map
$\mathrm{colim}_{{n}\geq 1}{\mathcal{Q}}^{\vee}\circ_{{\mathrm{f}}_{n}({\mathcal{Q}}^{\vee})}{\mathrm{A}}\to{\mathrm{A}}$
[…]
$\mathrm{colim}_{{n}\geq 1}{\mathrm{triv}}\circ_{{\mathrm{f}}_{n}({\mathcal{Q}}^{\vee})}{\mathrm{A}}\simeq{\mathrm{triv}}\circ_{{\mathcal{Q}}^{\vee}}(\mathrm{colim}_{{n}\geq 1}{\mathcal{Q}}^{\vee}\circ_{{\mathrm{f}}_{n}({\mathcal{Q}}^{\vee})}{\mathrm{A}})\to{\mathrm{triv}}\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}$
[…]
$\theta_{\mathcal{C}}:(\tau_{n}({\mathcal{Q}})\circ^{{\mathcal{Q}}}{\mathrm{triv}})\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}\simeq({\mathrm{triv}}\circ_{{\mathrm{f}}_{n}({\mathcal{Q}}^{\vee})}{\mathcal{Q}}^{\vee})\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}\to{\mathrm{triv}}\circ_{{\mathrm{f}}_{n}({\mathcal{Q}}^{\vee})}{\mathrm{A}}$
@@
Now let $\mathcal{C}$} be an arbitrary stable symmetric monoidal $\infty$-category that admits small colimits and totalizations. By Corollary 5.11 there are symmetric monoidal exact embeddings $\iota:{\mathcal{C}}\subset{\mathcal{C}}^{\prime},{j}:{\mathcal{C}}^{\prime}\subset{\mathcal{C}}^{\prime\prime}$ into stable symmetric monoidal $\infty$-categories, where $\mathcal{C}$$^{\prime\prime}$ is a presentably symmetric monoidal $\infty$-category, $\mathcal{C}$$^{\prime}$ is compatible with small colimits, the first embedding $\iota:{\mathcal{C}}\subset{\mathcal{C}}^{\prime}$ admits a left adjoint $\mathrm{L}$}, the second embedding ${j}:{\mathcal{C}}^{\prime}\subset{\mathcal{C}}^{\prime\prime}$ preserves small colimits and the whole embedding ${j}\circ\iota:{\mathcal{C}}\subset{\mathcal{C}}^{\prime\prime}$ preserves limits. We obtain induced embeddings
$\iota:{\mathrm{coOp}}({\mathcal{C}})\subset{\mathrm{coOp}}({\mathcal{C}}^{\prime}),{j}:{\mathrm{coOp}}({\mathcal{C}}^{\prime})\subset{\mathrm{coOp}}({\mathcal{C}}^{\prime\prime}),$
$\iota:\mathrm{Alg}_{{\mathcal{Q}}^{\vee}}({\mathcal{C}})\subset\mathrm{Alg}_{{\mathcal{Q}}^{\vee}}({\mathcal{C}}^{\prime}),{j}:\mathrm{Alg}_{{\mathcal{Q}}^{\vee}}({\mathcal{C}}^{\prime})\subset\subset\mathrm{Alg}_{{\mathcal{Q}}^{\vee}}({\mathcal{C}}^{\prime\prime})$
@@
${j}:{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1})\subset{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime\prime})_{\geq 1}),$
$\iota:\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\subset\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1}),$
${j}:\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime})_{\geq 1})\subset\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime\prime})_{\geq 1}).$
Since the embedding ${j}\circ\iota:{\mathcal{C}}\subset{\mathcal{C}}^{\prime\prime}$ preserves limits, the functor
$(-)\circ^{{j}\iota({\mathcal{Q}})}{j}\iota({\mathrm{triv}})\simeq(-)\ast^{{j}\iota({\mathcal{Q}})}{j}\iota({\mathrm{triv}}):\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime\prime})_{\geq 1})\to{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}}^{\prime\prime})_{\geq 1})$
@@
$(-)\circ^{{\mathcal{Q}}}{\mathrm{triv}}\simeq(-)*^{{\mathcal{Q}}}{\mathrm{triv}}:\mathrm{co}{\mathrm{RMod}}_{{\mathcal{Q}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1})\to{\mathrm{RMod}}_{{\mathcal{Q}}^{\vee}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}).$
Let ${\mathrm{T}}:=\tau_{n}({\mathcal{Q}})\circ^{{\mathcal{Q}}}{\mathrm{triv}}$. Consequently, there is a canonical equivalence:
${j}\iota({\mathrm{T}})\simeq{j}\iota(\tau_{n}({\mathcal{Q}}))\circ^{{j}\iota({\mathcal{Q}})}{j}\iota({\mathrm{triv}})\simeq\tau_{n}({j}\iota({\mathcal{Q}}))\circ^{{j}\iota({\mathcal{Q}})}{j}\iota({\mathrm{triv}}).$
@@
${j}\iota({\mathrm{triv}})\circ_{{j}\iota({\mathcal{Q}})^{\vee}}{j}\iota({\mathrm{A}})\simeq\mathrm{colim}_{{n}\geq 1}((\tau_{n}({j}\iota({\mathcal{Q}}))\circ^{{j}\iota({\mathcal{Q}})}{j}\iota({\mathrm{triv}}))\circ_{{j}\iota({\mathcal{Q}})^{\vee}}{j}\iota({\mathrm{A}}))$
$\simeq\mathrm{colim}_{{n}\geq 1}({j}\iota({\mathrm{T}})\circ_{{j}\iota({\mathcal{Q}})^{\vee}}{j}\iota({\mathrm{A}})).$
Since the embedding ${j}:{\mathcal{C}}^{\prime}\subset{\mathcal{C}}^{\prime\prime}$ preserves small colimits, there are canonical equivalences
${j}\iota({\mathrm{triv}})\circ_{{j}\iota({\mathcal{Q}})^{\vee}}{j}\iota({\mathrm{A}})\simeq{j}\iota({\mathrm{triv}})\circ_{{j}\iota({\mathcal{Q}}^{\vee})}{j}\iota({\mathrm{A}})\simeq{j}(\iota({\mathrm{triv}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})),$
${j}\iota({\mathrm{T}})\circ_{{j}\iota({\mathcal{Q}})^{\vee}}{j}\iota({\mathrm{A}})\simeq{j}\iota({\mathrm{T}})\circ_{{j}\iota({\mathcal{Q}}^{\vee})}{j}\iota({\mathrm{A}})\simeq{j}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})).$
@@
${j}(\iota({\mathrm{triv}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))\simeq\mathrm{colim}_{{n}\geq 1}{j}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))$
$\simeq{j}(\mathrm{colim}_{{n}\geq 1}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})))$
@@
$\iota({\mathrm{triv}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})\simeq\mathrm{colim}_{{n}\geq 1}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))$
@@
${\mathrm{triv}}\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}\simeq{\mathrm{L}}(\iota({\mathrm{triv}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))\simeq{\mathrm{L}}(\mathrm{colim}_{{n}\geq 1}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}})))\simeq$
$\mathrm{colim}_{{n}\geq 1}{\mathrm{L}}(\iota({\mathrm{T}})\circ_{\iota({\mathcal{Q}}^{\vee})}\iota({\mathrm{A}}))\simeq\mathrm{colim}_{{n}\geq 1}({\mathrm{T}}\circ_{{\mathcal{Q}}^{\vee}}{\mathrm{A}}).$
@@
Let $\mathcal{C}$} be a stable symmetric monoidal $\infty$-category that admits small limits, $\mathcal{O}$} a non-unital $\infty$-operad in $\mathcal{C}$} such that ${\mathcal{O}}_{1}={\mathbb{1}}$. For every ${\mathrm{C}}\in\mathrm{co}\mathrm{Alg}^{\mathrm{dp}}_{{\mathcal{O}}^{\vee}}({\mathcal{C}})$ in $\mathcal{C}$} there is a canonical equivalence in $\mathcal{C}$}:
${\mathrm{triv}}*^{{\mathcal{O}}^{\vee}}{\mathrm{C}}\simeq\lim_{{n}\geq 1}((\tau_{n}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})*^{{\mathcal{O}}^{\vee}}{\mathrm{C}}).$